The DT group is undertaking substantial work both for detector maintenance and for detec-tor upgrade.
Maintenance interventions on chambers and minicrates require close collaboration between DT, RPC and HO, and are difficult because they depend on the removal of thermal shields and cables on the front and rear of the chambers in order to gain access. The tasks are particularly critical on the central wheel due to the presence of fixed services. Several interventions on the chambers require extraction of the DT+RPC package: a delicate operation due to the very limited space for handling the big chambers, and the most dangerous part of the DT maintenance campaign. The interventions started in July 2013 and will go on until spring 2014. So far out of the 16 chambers with HV problems, 13 have been already repaired, with a global yield of 217 recovered channels. Most of the observed problems were due to displacement of impurities inside the gaseous volume. For the minicrates and FE, repairs occurred on 22 chambers in YB–1 and 14 in YB–2. Maintenance of the YB–1 and YB–2 wheels has finished.

Upgrade activities continue to evolve in good accordance with the schedule, both for the theta Trigger Board (TTRB) replacement and for the Sector Collector (SC) relocation from UXC to USC:

  • The TTRB work aims at reconstituting the stock of spare boards for the long-term op-eration of the chamber minicrates. The intervention has been done on all minicrates of MB1 chambers in wheel YB–2, and on some YB+2 MB1 minicrates.
  • With the SC work, data and trigger primitives from each of the 250 DT chambers will be available in USC on optical fibers. This work is the corner stone for any long-term upgrade plan of the DT system. All optical-fiber trunk cables have been received, verified and installed between USC and UXC. All of the Wiener crates are installed.
 Pro-duction of all  CUOF mezzanines, OFCU-RO boards and OFCU-TRG boards has finished: the success rate for passing QC tests and burn-in is very high.
 CUOF mezzanines have been assembled on CUOF motherboards.
  • SC relocation has been performed for wheel YB–1, and the new electronics cabled up. System commissioning is going on with test pulser and cosmic data: first measurements of the new trigger latency show an increase of +3BX as expected by design.

Regarding Phase 2 upgrade strategies, design studies have started for rebuilding the minicrates, in view of their installation during LS3. The studies are also revisiting the entire architecture of DT back-end RO and trigger electronics, aiming at simplification.

by Marco Dallavalle